# Optimal and Practical Batched Linear Bandit Algorithm

### v1

@@ 1 Introduction
Batched linear bandit algorithms must strike a delicate balance between exploration and exploitation within each batch, aiming to minimize both the number of policy updates and the cumulative regret. Although recent work provides strong theoretical guarantees (Ren et al., 2024; Zhang et al., 2025), their empiricial limitations may undermine the intended benefits of batched learning. In numerical experiments (see Section 7), the method of Ren et al. (2024) tends to eliminate arms prematurely (often including the optimal arm) and suffers in practical performance, while other methods update parameters rather too frequently (Abbasi-Yadkori et al., 2011; Esfandiari et al., 2021). [^1] Note that Abbasi-Yadkori et al. (2011, Section 5.1) introduced a rare-switching variant of OFUL that attains a batch complexity of $\mathcal{O}(d\log T)$. Still others require such substantial computation that they become infeasible even at moderate problem scales (Ruan et al., 2021; Hanna et al., 2023; Zhang et al., 2025) (see Appendix E). Much of this empirical inefficiency observed with the methods in Ruan et al. (2021); Hanna et al. (2023); Zhang et al. (2025) stems from algorithmic components tailored to contextual settings with time-varying features; when these algorithms are applied to the plain, fixed-feature linear bandit, they cannot exploit the static structure and therefore still incur excessive computation.
@@
  • Novel Batched Algorithm and Regret Guarantees. We propose BLAE, which constructs a judiciously designed arm-elimination strategy integrated with regularized G-optimal design. BLAE achieves a worst-case regret (Theorem 1) of
@@
  • Refined Technical Analysis. Key to our results are new analysis techniques for regularized G-optimal design under batch updates (Lemma 3), along with a concentration bounds that accommodate both batching and regularization (Lemmas 1 and 2). Additionally, our refined analysis of efficient coverings of the unit sphere provides both theoretical guarantees and practical improvements (Lemma 8). These technical tools may be of independent interest for both batched and non-batched bandit research.
@@
Zhang et al. (2025) $\mathcal{O}(\sqrt{dT\log K}\,\text{poly}\log(dT)\log\log T)^{\dagger}$ $\mathcal{O}(\log\log T)$
@@
  • $\dagger$
    The term $\text{poly}\log(dT)$ represents a polynomial function of $\log(dT)$ with a degree of at least one.
@@ 2 Related Work
Despite these theoretical advances, many existing methods face practical obstacles. Algorithms that rely on $\mathcal{O}(\log T)$ batches (Abbasi-Yadkori et al., 2011; Esfandiari et al., 2021) still update more frequently than the desirable $\mathcal{O}(\log\log T)$, increasing computational overhead. Approaches that do match the lower batch complexity often employ two-phase updates that split each batch into separate “estimation” and “policy-learning” stages (Ruan et al., 2021; Zhang et al., 2025); this structure inflates runtime and may leave the information matrix ill-conditioned. Furthermore, the algorithms of Ruan et al. (2021), Hanna et al. (2023), and Zhang et al. (2025) incorporate subroutines tailored to time-varying features. Even after these routines are trimmed to suit our fixed-feature setting for comparison, the implementations of these methods still remain computationally intensive (see Appendix E), rendering the algorithms impractical even at moderate problem sizes. While it is fair to acknowledge that these subroutines were devised to handle more general, time-varying contexts, there remains no straightforward way to streamline them for the plain, fixed-feature linear bandit, leaving the computational burden largely unresolved.
A recent work (Ren et al., 2024) shows that $\mathcal{O}(\log\log T)$ batches suffice to achieve the regret $\widetilde{\mathcal{O}}\!\bigl{(}\sqrt{dT\log K}\bigr{)}$ in the small-$K$ regime ($\log K\leq\mathcal{O}(d)$), the tightest bound known so far. Optimality in the large-$K$ regime ($\log K\geq\Omega(d)$) is still open, and empirical experiments (Section 7) reveal notable weaknesses: the algorithm often eliminates arms prematurely—including the optimal arm—and is highly sensitive to the hyperparameter $\gamma$, whose choice can enlarge the third batch to $\mathcal{O}\!\bigl{(}(\log T)^{1+\gamma}\bigr{)}$, exhausting the horizon and impeding exploitation.
@@ 3.1 Notations
We denote the cardinality of a set by ||$\cdot$|| ⋅ |, and the indicator function by $\text{1}_{A}(x)$, where $1\text{1}_{A}(x)=$ 1 if $x\in A$, and $0\text{1}_{A}(x)=$ 0 otherwise. For a vector $x\in\mathbb{R}^{d}$, the $\ell_{2}$-norm is denoted by $\|x\|$, and the weighted $\ell_{2}$-norm associated with a positive definite matrix $H$ is defined as $\|x\|_{H}=\sqrt{x^{T}Hx}.$ The trace, determinant, and adjugate (the transpose of the cofactor matrix) of $H$ are denoted by $\text{trace}(H)$, $\det(H)$, and $\text{adj}(H)$ respectively. We denote by $[n]$ the set $\{1,\ldots,n\}$ for a positive integer $n$. We denote $\wedge$ as the minimum operator.
@@ 3.2 Problem Setting: Batched Linear Bandits
where $t_{0}=0$, $t_{B}=T$, and $B$ is the total number of allowed policy updates. A batch is thus a contiguous block of rounds during which the agent must use a fixed policy. In other words, within each batch, the agent is not able to update its policy, effectively deferring feedback from that entire batch until the end. Only at the end of batch $\ell$ can the agent update its policy using the rewards collected from rounds $t_{\ell-1}+1$ to $t_{\ell}$. Consequently, the agent has at most $B-1$ opportunities to adapt its strategy over the entire horizon, rather than $T-1$ as in the fully adaptive setting.
@@ 4 Algorithm: BLAE
We introduce Batched Linear bandit with Arm Elimination (BLAE), whose pseudocode is provided in Algorithm 1. As its name suggests, our algorithm BLAE sequentially eliminates arms and updates the available arm set $\mathcal{A}_{\ell-1}=\{x^{(1)},\ldots,x^{(|\mathcal{A}_{\ell-1}|)}\}$ after completing each batch $(\ell-1)$.
@@
$\mathcal{D_{\ell}}\coloneqq\bigg{\{}\pi_{\ell}\in\mathbb{R}^{|\mathcal{A}_{\ell-1}|}\,\bigg{|}\,\sum_{i=1}^{|\mathcal{A}_{\ell-1}|}\pi_{\ell i}=1\bigg{\}}\;.$
@@
Here, while the construction of the elimination threshold $\varepsilon_{\ell}$ is crucial for our theoretical analysis (its order is specified in Theorem 1 and the actual values are specified in Lemma 2), to maximize practical performance, we use the bounds without simplification, resulting in complex relational expressions rather than straightforward forms.
@@
8:        Pull arm $x^{(i)}$ repeatedly for the following times$\Big{\{}\lceil T^{\frac{2^{\ell}-1}{2^{\ell}}}\!\!\left(c_{\ell}\pi^{*}_{\ell i}+(1\!-\!c_{\ell})\!\cdot\!\text{1}_{\{1\}}\!(i)\right)\rceil\,\wedge\,T\!-\!t\Big{\}}$;
@@ 5 Regret Analysis
As we delve into the regret analysis of our proposed algorithm (Section 4), we first present the standard assumptions used in the literature.
@@ 5.1 Assumptions
The noise $\eta_{t}$ is a 1-subgaussian random variable for all time steps $t=1,...,T$.
@@ 5.2 Worst-Case Regret of BLAE
Discussion of Theorem 1. Theorem 1 shows that BLAE achieves a regret bound of $\widetilde{\mathcal{O}}\bigl{(}\sqrt{dT\log K}\wedge d\sqrt{T}\bigr{)}$, matching the best-known (and indeed minimax-optimal) results for fully adaptive linear bandit problems. Moreover, BLAE operates with only $\mathcal{O}(\log\log T)$ batches, comparable to the lowest batch complexity among existing batched linear bandit methods. Critically, this bound simultaneously covers both the large-$K$ regime ($K\geq\Omega(e^{d})$) with $\mathcal{O}(d\sqrt{T})$ regret and the small-$K$ regime ($K\leq\mathcal{O}(e^{d})$) with $\mathcal{O}(\sqrt{dT\log K})$ regret, thereby establishing the tightest known performance guarantees for batched linear bandits in both regimes. To our knowledge, this is the first time such matching lower bounds have been simultaneously achieved under limited adaptivity.
Central to our result are new techniques for regularized G-optimal design under batch updates (see Lemma 3), alongside a rigorous concentration bound that handles both batching and regularization (Lemmas 1 and 2). We also refine the construction of unit-sphere coverings to optimize batch-wise exploration (Lemma 8). These technical contributions may be of independent interest for other batched (and potentially non-batched) bandit problems.
@@ 5.3 Proof Sketch of Theorem 1
After the $\ell$-th batch $(\ell\geq 1)$, the following results holds. For any $x\in\mathbb{R}^{d}$ and $10<\delta<$ 1
@@
Lemma 1 establishes that the inner product between the estimation error and an arbitrary vector $x$ is upper bounded by the Mahalanobis norm $\|x\|_{H_{\ell}^{-1}}$ with high probability.
@@
Lemma 2 provides a high probability guarantee that the estimation error remains tightly controlled across all batches and arm pairs. The term $\beta_{\ell 1}$ quantifies the difference between the batched regularized least squares estimator $\hat{\theta}_{\ell}$ and the unknown parameter $\theta^{*}$, which plays a crucial role in establishing the regret bound of $\widetilde{\mathcal{O}}(d\sqrt{T})$ in the large-$K$ regime. Similarly, $\beta_{\ell 2}$ captures the discrepancy between $\hat{\theta}_{\ell}$ and $\theta^{*}$ from a different perspective, which is essential for deriving the regret bound of $\widetilde{\mathcal{O}}(\sqrt{dT\log K})$ in the small-$K$ regime. By defining $\beta_{\ell}(\delta)\coloneqq\beta_{\ell 1}(\delta)\wedge\beta_{\ell 2}(\delta)$ as the minimum of these two terms, our framework seamlessly handles both regimes. This motivates the construction of $\varepsilon_{\ell}(\delta)$ as defined above. For convenience, we define
@@
Lemma 3 establishes an upper bound on the uncertainty, measured by the Mahalanobis norm $\|x\|_{H_{\ell}^{-1}}^{2}$, for any arm in the $\ell$-th batch. Most G-optimal design methods used in bandit algorithms omit regularization and rely solely on the Kiefer-Wolfowitz equivalence theorem (Kiefer & Wolfowitz, 1960). However, similar to the argument in Lemma 1, under an arm elimination setting, the information matrix $H_{\ell}$ may not remain invertible without regularization. Thus, a refined analysis that explicitly incorporates the regularization term is necessary. This lemma further demonstrates that we can derive an even tighter upper bound on the uncertainty compared to approaches that do not account for regularization. The exponential decay term $T^{-\frac{2^{\ell}-1}{2^{\ell}}}$ ensures that uncertainty decreases rapidly as the batch index $\ell$ increases and the time horizon $T$ grows, highlighting the algorithm’s adaptability to batch-wise updates. By leveraging the bound $\Delta_{x}\leq 2\varepsilon_{\ell-1}(\delta)$, which holds for all arms available in the $\ell$-th batch with high probability, and applying Lemma 3, we can effectively bound the expected cumulative regret as described in Theorem 1.
@@ 6 Why Many Existing Provable Batched Linear Bandit Algorithms Underperform in Practice?
Other algorithms also encounter significant challenges in practical evaluations of batched linear bandits. For instance, BatchLinUCB-DG (Ruan et al., 2021) splits each batch into two phases: the second phase depends on newly gathered context to refine the policy, a requirement that proves ineffectual when arms are fixed across rounds. Our numerical experiments consistently show that BatchLinUCB-DG struggles to execute meaningful exploration, frequently failing to “double” the determinant of its information matrix and thus settling on degenerate (often zero) exploration distributions. Additionally, the algorithm’s repeated arm elimination via a UCB rule and subsequent mixed-softmax policy learning imposes high computational overhead, rendering it impractical at larger scales.
@@
Lastly, Algorithm 1 in Zhang et al. (2025) adopts a similar two-phase batch-splitting approach as in Ruan et al. (2021), thereby inheriting the similar limitations. As with its counterpart, the algorithm yields suboptimal policies in fixed-arm environments and involves considerable computational overhead at every batch update. In particular, it leans heavily on re-estimating parameters and allocating exploratory actions within each batch, which becomes exceedingly costly in computation when the action set is large or the time horizon $T$ grows. Our runtime experiments in Table 2 confirm that these approaches (Ruan et al., 2021; Hanna et al., 2023; Zhang et al., 2025) struggle to scale efficiently even for moderately sized $T$ and other problem parameters, rendering them impractical for large problem instances. More broadly, the combination of premature arm elimination, extensive computation at each batch, and lack of consideration for independence requirements in batched settings explains why methods that appear sound theoretically may struggle to deliver in real-world applications.

### v2

@@ 1 Introduction
Batched linear bandit algorithms must strike a delicate balance between exploration and exploitation within each batch, aiming to minimize both the number of policy updates and the cumulative regret. Although recent work provides strong theoretical guarantees (Ren et al., 2024; Zhang et al., 2025), their empiricial limitations may undermine the intended benefits of batched learning. In numerical experiments (see Section˜7), the method of Ren et al. (2024) tends to eliminate arms prematurely (often including the optimal arm) and suffers in practical performance, while other methods update parameters rather too frequently (Abbasi-Yadkori et al., 2011; Esfandiari et al., 2021). [^1] Note that Abbasi-Yadkori et al. (2011, Section 5.1) introduced the rare-switching OFUL that attains a batch complexity of $\mathcal{O}(d\log T)$. Still others require such substantial computation that they become infeasible even at moderate problem scales (Ruan et al., 2021; Hanna et al., 2023; Zhang et al., 2025) (see Appendix˜E). Much of this empirical inefficiency observed with the methods in Ruan et al. (2021); Hanna et al. (2023); Zhang et al. (2025) stems from algorithmic components tailored to contextual settings with time-varying features; when these algorithms are applied to the plain, fixed-feature linear bandit, they cannot exploit the static structure and therefore still incur excessive computation.
@@
  • Novel Batched Algorithm and Regret Guarantees. We propose BLAE, which constructs a judiciously designed arm-elimination strategy integrated with regularized G-optimal design. BLAE achieves a worst-case regret (Theorem˜1) of
@@
  • Refined Technical Analysis. Key to our results are new analysis techniques for regularized G-optimal design under batch updates (Lemma˜3), along with a concentration bounds that accommodate both batching and regularization (Lemmas 1 and 2). Additionally, our refined analysis of efficient coverings of the unit sphere provides both theoretical guarantees and practical improvements (Lemma˜8). These technical tools may be of independent interest for both batched and non-batched bandit research.
@@
Zhang et al. (2025) $\mathcal{O}(\sqrt{dT\log(dKT)\log T}\,\log(dT)\log\log T)$ $\mathcal{O}(\log\log T)$
@@ 2 Related Work
Despite these theoretical advances, many existing methods face practical obstacles. Algorithms that rely on $\mathcal{O}(\log T)$ batches (Abbasi-Yadkori et al., 2011; Esfandiari et al., 2021) still update more frequently than the desirable $\mathcal{O}(\log\log T)$, increasing computational overhead. Approaches that do match the lower batch complexity often employ two-phase updates that split each batch into separate “estimation” and “policy-learning” stages (Ruan et al., 2021; Zhang et al., 2025); this structure inflates runtime and may leave the information matrix ill-conditioned. Furthermore, the algorithms of Ruan et al. (2021), Hanna et al. (2023), and Zhang et al. (2025) incorporate subroutines tailored to time-varying features. Even after pruning these components to fit our fixed-feature setting, the implementations remain computationally intensive and become impractical at moderate problem sizes (see Appendix E). Although those routines are justified in broader settings, we find no simple way to streamline them for the plain fixed-feature linear bandit, leaving much of the burden unresolved.
A recent work (Ren et al., 2024) shows that $\mathcal{O}(\log\log T)$ batches suffice to achieve the regret $\widetilde{\mathcal{O}}\!\bigl{(}\sqrt{dT\log K}\bigr{)}$ in the small-$K$ regime ($\log K\leq\mathcal{O}(d)$), the tightest bound known so far. However, designing an algorithm that achieves optimal regret in both the large-$K$ and small-$K$ regimes simultaneously remains an open problem. Empirical experiments (Section˜7) further reveal notable weaknesses: the algorithm often eliminates arms prematurely—including the optimal arm—and is highly sensitive to the hyperparameter $\gamma$, whose choice can enlarge the third batch to $\mathcal{O}\!\bigl{(}(\log T)^{1+\gamma}\bigr{)}$, exhausting the horizon and impeding exploitation.
@@ 3.1 Notations
We denote the cardinality of a set by $|\cdot|$, and the indicator function by $\text{1}_{A}(x)$, where $\text{1}_{A}(x)=1$ if $x\in A$, and $\text{1}_{A}(x)=0$ otherwise. For a vector $x\in\mathbb{R}^{d}$, the $\ell_{2}$-norm is denoted by $\|x\|_{2}$, and the weighted $\ell_{2}$-norm associated with a positive definite matrix $H$ is defined as $\|x\|_{H}=\sqrt{x^{T}Hx}.$ The trace, determinant, and adjugate (the transpose of the cofactor matrix) of $H$ are denoted by $\text{trace}(H)$, $\det(H)$, and $\text{adj}(H)$ respectively. We denote by $[n]$ the set $\{1,\ldots,n\}$ for a positive integer $n$, and use $\wedge$ to denote the minimum operator.
@@ 3.2 Problem Setting: Batched Linear Bandits
where $t_{0}=0$, $t_{B}=T$, and $B$ is the total number of allowed policy updates. A batch is thus a contiguous block of rounds during which the agent must use a fixed policy. In other words, within each batch, the agent is not able to update its policy, effectively deferring feedback from that entire batch until the end. Only at the end of batch $\ell$ can the agent update its policy using the rewards collected from rounds $t_{\ell-1}+1$ to $t_{\ell}$. Consequently, the agent has at most B1B-1 opportunities to adapt its strategy over the entire horizon, rather than T1T-1 as in the fully adaptive setting.
@@ 4 Algorithm: BLAE
We introduce Batched Linear bandit Algorithm with Elimination (BLAE), whose pseudocode is provided in Algorithm 1. As its name suggests, our algorithm BLAE sequentially eliminates arms and updates the available arm set $\mathcal{A}_{\ell}=\{x^{(1)},\ldots,x^{(|\mathcal{A}_{\ell}|)}\}$ after completing each batch $\ell$.
@@
$\mathcal{D_{\ell}}\coloneqq\bigg{\{}\pi_{\ell}\in\mathbb{R}_{\geq 0}^{|\mathcal{A}_{\ell-1}|}\,\bigg{|}\,\sum_{i=1}^{|\mathcal{A}_{\ell-1}|}\pi_{\ell i}=1\bigg{\}}\;.$
@@
Here, while the construction of the elimination threshold $\varepsilon_{\ell}$ is crucial for our theoretical analysis (its order is specified in Theorem˜1 and the actual values are specified in Lemma 2), to maximize practical performance, we use the bounds without simplification, resulting in complex relational expressions rather than straightforward forms.
@@
[…]
@@ 5 Regret Analysis
As we delve into the regret analysis of our proposed algorithm (Section˜4), we first present the standard assumptions used in the literature.
@@ 5.1 Assumptions
The noise $\eta_{t}$ is a 1-subgaussian random variable for all time steps $t\in[T]$.
@@ 5.2 Worst-Case Regret of BLAE
Discussion of Theorem˜1. Theorem 1 shows that BLAE achieves a regret bound of $\widetilde{\mathcal{O}}\bigl{(}\sqrt{dT\log K}\wedge d\sqrt{T}\bigr{)}$, matching the best-known (and indeed minimax-optimal) results for fully adaptive linear bandit problems. Moreover, BLAE operates with only $\mathcal{O}(\log\log T)$ batches, comparable to the lowest batch complexity among existing batched linear bandit methods. Critically, this bound simultaneously covers both the large-$K$ regime ($K\geq\Omega(e^{d})$) with $\mathcal{O}(d\sqrt{T})$ regret and the small-$K$ regime ($K\leq\mathcal{O}(e^{d})$) with $\mathcal{O}(\sqrt{dT\log K})$ regret, thereby establishing the tightest known performance guarantees for batched linear bandits in both regimes. To our knowledge, this is the first time such matching lower bounds have been simultaneously achieved under limited adaptivity.
Central to our result are new techniques for regularized G-optimal design under batch updates (see Lemma˜3), alongside a rigorous concentration bound that handles both batching and regularization (Lemmas 1 and 2). We also refine the construction of unit-sphere coverings to optimize batch-wise exploration (Lemma˜8). These technical contributions may be of independent interest for other batched (and potentially non-batched) bandit problems.
@@ 5.3 Proof Sketch of Theorem 1
After the $\ell$-th batch $(\ell\geq 1)$, the following results holds. For any $x\in\mathbb{R}^{d}$ and $0<\delta<1$
@@
Lemma˜1 establishes that the inner product between the estimation error and an arbitrary vector $x$ is upper bounded by the Mahalanobis norm $\|x\|_{H_{\ell}^{-1}}$ with high probability.
@@
Lemma˜2 provides a high probability guarantee that the estimation error remains tightly controlled across all batches and arm pairs. The term $\beta_{\ell 1}$ quantifies the difference between the batched regularized least squares estimator $\hat{\theta}_{\ell}$ and the unknown parameter $\theta^{*}$, which plays a crucial role in establishing the regret bound of $\widetilde{\mathcal{O}}(d\sqrt{T})$ in the large-$K$ regime. Similarly, $\beta_{\ell 2}$ captures the discrepancy between $\hat{\theta}_{\ell}$ and $\theta^{*}$ from a different perspective, which is essential for deriving the regret bound of $\widetilde{\mathcal{O}}(\sqrt{dT\log K})$ in the small-$K$ regime. By defining $\beta_{\ell}(\delta)\coloneqq\beta_{\ell 1}(\delta)\wedge\beta_{\ell 2}(\delta)$ as the minimum of these two terms, our framework seamlessly handles both regimes. This motivates the construction of $\varepsilon_{\ell}(\delta)$ as defined above. For convenience, we define
@@
Lemma˜3 establishes an upper bound on the uncertainty, measured by the Mahalanobis norm $\|x\|_{H_{\ell}^{-1}}^{2}$, for any arm in the $\ell$-th batch. Most G-optimal design methods used in bandit algorithms omit regularization and rely solely on the Kiefer-Wolfowitz equivalence theorem (Kiefer & Wolfowitz, 1960). However, similar to the argument in Lemma 1, under an arm elimination setting, the information matrix $H_{\ell}$ may not remain invertible without regularization. Thus, a refined analysis that explicitly incorporates the regularization term is necessary. This lemma further demonstrates that we can derive an even tighter upper bound on the uncertainty compared to approaches that do not account for regularization. The exponential decay term $T^{-\frac{2^{\ell}-1}{2^{\ell}}}$ ensures that uncertainty decreases rapidly as the batch index $\ell$ increases and the time horizon $T$ grows, highlighting the algorithm’s adaptability to batch-wise updates. By leveraging the bound $\Delta_{x}\leq 2\varepsilon_{\ell-1}(\delta)$, which holds for all arms available in the $\ell$-th batch with high probability, and applying Lemma˜3, we can effectively bound the expected cumulative regret as described in Theorem˜1.
@@ 6 Why Many Existing Provable Batched Linear Bandit Algorithms Underperform in Practice?
Other algorithms also encounter significant challenges in practical evaluations of batched linear bandits. For instance, BatchLinUCB-DG (Ruan et al., 2021) splits each batch into two phases: the second phase depends on newly gathered context to refine the policy, a requirement that proves ineffectual when arms are fixed across rounds. Our numerical experiments consistently show that BatchLinUCB-DG struggles to execute meaningful exploration, frequently failing to “double” the determinant of its information matrix and thus settling on degenerate (often zero) exploration distributions. In addition, its mixed-softmax policy, which incorporates a G-optimal design at every round, incurs substantial computational overhead, thereby restricting its applicability to larger-scale problems.
@@
Lastly, Algorithm 1 in Zhang et al. (2025) adopts a two-phase batch-splitting strategy similar to that of Ruan et al. (2021), thereby inheriting the similar limitations. As with its counterpart, the algorithm produces suboptimal policies in fixed-arm environments and incurs considerable computational overhead from computing a G-optimal design at every round of the first batch, a cost that grows rapidly when the action set is large or the time horizon $T$ increases.
Our runtime experiments in Table 2 confirm that these approaches (Ruan et al., 2021; Hanna et al., 2023; Zhang et al., 2025) struggle to scale efficiently even for moderately sized $T$ and other problem parameters, rendering them impractical for large-scale problem instances. Moreover, all three methods (Ruan et al., 2021; Hanna et al., 2023; Zhang et al., 2025) assume that contexts are drawn independently and identically distributed (i.i.d.). In fixed-arm environments, the contexts remain identically distributed but are not independent, violating this assumption and undermining the theoretical validity of these methods in our setting.
More broadly, the combination of premature arm elimination, extensive computational costs, and the lack of consideration for independence requirements in batched settings helps explain why methods that appear theoretically sound may struggle to deliver in real-world applications.